# Perfect supercurrent diode efficiency in chiral nanotube-based Josephson junctions

### v1

@@
Abstract
@@
The supercurrent diode effect (SDE) describes superconducting systems where the magnitude of the superconducting-to-normal state switching current differs for positive and negative current bias. Despite the ubiquity of such diode effects in Josephson devices, the fundamental conditions to observe a diode effect in a Josephson junction and achieve perfect diode efficiency remain unclear. In this work, we analyze the supercurrent diode properties of a chiral nanotube-based Josephson junction within a Ginzburg-Landau theory. We find a diode effect and anomalous phase develops across the junction when a magnetic field is applied parallel to the tube despite the absence of spin-orbit interactions in the system. Unexpectedly, the SDE in the junction is independent of the anomalous phase. Alternatively, we determine a non-reciprocal persistent current that is protected by fluxoid quantization can activate SDE, even in the absence of higher-order pair tunneling processes. We show this new type of SDE can lead to, in principle, a perfect diode efficiency, highlighting how persistent currents can be used to engineer high efficiency supercurrent diodes.
@@
With a growing need for low-power fast electronics with low operational temperatures, a rapid development of non-reciprocal superconducting devices has occurred recently 1, 2. Non-reciprocal effects have been intensely investigated in bulk superconductors 3, 4, 5, 6, 7 as well as in Josephson junctions 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19 (JJs)– two superconducting electrodes weakly coupled by a tunneling barrier, such as a normal metal or insulator. In JJs, the Josephson diode effect (JDE) manifests as a difference in magnitude of positive and negative threshold currents where the device switches from a superconducting to a dissipative state, shown schematically in Fig. 1(a). Despite the effect having been known decades ago in devices with geometric inhomogeneities, the subject has received renewed interest based on superconducting diode effects reported in homogeneous devices where non-reciprocity arises from microscopic interactions. In these systems the diode effect, particularly when no magnetic field is applied, has supported the discovery of exotic states, such as time-reversal symmetry broken superconducting states 20, 21. Other zero-field supercurrent diodes include JJs made of iron-based superconductors 22, JJs of twisted bilayer graphene 23, twisted trilayer graphene 24, obstructed atomic insulator JJs 25, strained PbTaSe2 26, and multiferroic JJs 27. Open questions linger about the nature of the SDE in some of these systems and others, which calls for further theoretical modeling to address the fundamental limits of the SDE.
@@
Figure 1: (a) Schematic of voltage-current curve of a supercurrent diode having a negative diode polarity. (b) Schematic of $\mathcal{C}_{2}$-symmetric Fermi surface. (c) Illustration of chiral nanotube on a square lattice. (d) Cartoon of the ChNt-JJs and the helical persistent current inducing the Josephson diode effect.
@@
Theoretical descriptions of the JDE often heuristically focus on symmetry arguments, namely, broken inversion and time-reversal symmetries 28, 29, 30, 1. Assuming these symmetries are broken, one may consider a common minimal expression for the current-phase relationship (CPR) 2: $I_{s}(\phi)=a\sin(\phi)+b\cos(\phi)+c\sin(2\phi)+d\cos(2\phi).$ Here $\phi$ is the phase across the Josephson junction, and $a,b,c$ and $d$ are real-valued constants describing the weights describing phase-coherent Cooper pair tunneling ($a$ and $b$) and pair co-tunneling ($c$ and $d$) supercurrent channels. The parameters $b$ and $d$ are associated with broken time-reversal symmetry (TRS) in the junction. When all four constants are non-zero and treated as independent parameters, then ubiquitous combinations of $(a,b,c,d)$ result in $I_{c+}\neq|I_{c-}|$, where $I_{c+}=\max\left(I_{s}\right)$ and $I_{c-}=\min\left(I_{s}\right),$ and the Josephson diode effect is realized (see Fig. 1(a)) with an efficiency $\eta=\frac{I_{c+}+I_{c-}}{I_{c+}-I_{c-}}$. Using the minimal form of $I_{s}(\phi)$, it is apparently not possible to have a perfectly efficient supercurrent diode where either $I_{c+}$ or $I_{c-}$ are zero. While ideal diode operation has been identified in the ac limit 31, 32, 33, there remains a question in the dc case: is it possible, in principle, to have a perfectly efficient Josephson diode operating in the dc limit? We answer this question in the affirmative by considering chiral nanotube-based Josephson junctions (ChNt-JJs).
@@
In this work, we present a phenomenological Ginzburg-Landau (GL) theory for a ChNt-JJ with a magnetic field applied along the tube. We consider an anisotropic free energy functional that obeys inversion symmetry and choose periodic boundary conditions along non-high symmetry axes to define a chiral nanotube, see Fig. 1(b-c). We first show that in this system, the ChNt-JJ develops an anomalous phase despite the absence of spin-orbit coupling. The link between the anomalous phase and JDE has been investigated in JJs that break inversion symmetry 34, but the existence and possible origin of an anomalous phase in ChNt-JJs has not been identified until now. We also find a diode effect in the absence of pair co-tunneling, due to a non-reciprocal persistent current which is protected by fluxoid quantization in the ChNt-JJ. In this case, a phase-independent persistent current allows us to place an upper bound on the diode efficiency demonstrating how perfect efficiency can be achieved in ChNt-JJs.
@@
We model a superconducting ChNt with higher-order terms in the free energy functional 35:
@@
Here $m_{0}<m_{1},m_{2}$, $\zeta_{0},\zeta_{1},\zeta_{2}>0$, $\alpha\propto(T-T_{c})$ and $\beta$ are the usual GL coefficients, and $\mathbf{p}=-i\hbar\left(\nabla-i\frac{2e}{\hbar c}\mathbf{A}\right)$ is the momentum operator. The atomic lattice of the system is defined on a 2D sheet $\Omega\in\mathbb{R}^{2}$. Here the kinetic energy contribution to $F$ has inversion symmetry and $\mathcal{C}_{2}$ rotational symmetry when $1/m_{1}\neq 0$ This corresponds to an elliptical Fermi surface, see Fig. 1(b). The circumferential vector $\mathbf{C}=2\pi R(-\sin\theta,\cos\theta),$ where $R$ is the radius of the nanotube, defines the periodic boundary conditions to apply to the nanotube, see Fig. 1(c). To simplify our analysis, we rotate our real space basis $\mathbf{r}=R(\theta)\mathbf{r}_{0}$ where $R(\theta)$ is the 2x2 rotation matrix. The Ginzburg-Landau equations in this case are (see Supplementary Information for details)
@@
$\left[\alpha+\beta{\left|\psi\right|}^{2}-\left(\rho_{1}p_{x}^{2}+\rho_{2}p_{y}^{2}+2\rho_{3}p_{x}p_{y}\right)+\rho_{4}\left(p_{x}^{4}+p_{y}^{4}\right)+\left(\rho_{6}+2\rho_{7}\right)p_{x}^{2}p_{y}^{2}+2\rho_{8}\left(\rho_{x}^{3}p_{y}-p_{x}p_{y}^{3}\right)\right]\psi=0,$ (3)
@@
To establish the conditions for the SDE, we consider an external magnetic field along the nanotube $\mathbf{B}=B_{ext}\mathbf{x}$ which breaks time-reversal symmetry. Periodic boundary conditions lead to $p_{y}$ being a good quantum number $p_{y}^{(n)}=\hbar(\pi R^{2}B_{ext}/\Phi_{0}+n)/R$ for some $n\in\mathbb{Z}$ and where $\Phi_{0}$ is the flux quantum. We restrict ourselves to the case of a small diameter nanotube and the lowest sub-band for simplicity, and henceforth set $p_{y}^{(n)}=p_{y}^{(0)}$. Then the order parameter is
@@
for some complex-valued function $\psi_{x}$. From here, to calculate the supercurrent in a Josephson junction, we need to apply appropriate boundary conditions to the solution to Eq. (3) in the junction. The simplest model for a Josephson junction uses rigid boundary conditions 36
$\psi_{x}(x\leq 0)=\psi_{\infty},\quad\psi_{x}(x\geq L)=\psi_{\infty}e^{i\phi},$ (5)
where the normal region defining the junction is $0\leq x\leq L$, $\phi$ is the phase difference across the junction, and $\psi_{\infty}=\sqrt{-\alpha/\beta}.$ A more rigorous treatment of the boundary conditions for the junction can be made and may be important in 2D and 3D geometries 37, but for our quasi-1D system rigid boundary conditions provide sufficient qualitative accuracy. Assuming $T\lesssim T_{c}$ and a short junction $L\ll\xi$ where $\xi$ is the superconducting coherence length, we can solve Eq. (3) by linearizing in the standard way and find
$\psi_{x}(x)$ $\approx\psi_{\infty}L^{-1}e^{iax}\left(L-x+xe^{i(\phi-aL)}\right),$ (6)
where $a=-\frac{\rho_{3}p_{y}^{(0)}}{\hbar\rho_{1}}$. Then the current-phase relationship of the ChNt-JJ is calculated by solving for $J_{x}$ in Eq. (2) which, due to the anisotropic superfluid stiffness, requires both components of $\bf j$:
$\bf{j}$ $=\begin{pmatrix}2i\frac{\psi_{\infty}^{2}}{L}\sin\left(\phi-aL\right)+2ia\psi_{0}^{2}{\left|L-x+x^{i(\phi-aL)}\right|}^{2}\\ \frac{2i\psi_{0}^{2}p_{y}^{(0)}}{\hbar}{\left|L-x+x^{i(\phi-aL)}\right|}^{2}\end{pmatrix}.$ (7)
@@
$I_{s}(\phi)$ $=I_{c}\left[\sin\tilde{\phi}+\frac{2\hat{\Phi}}{LR}\gamma^{-1}\left(1-\cos\tilde{\phi}\right)\right].$ (9)
Here $I_{c}=\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}$, $\tilde{\phi}=\phi-\phi_{0}$, $\phi_{0}=-\frac{\rho_{3}p_{y}^{(0)}L}{\hbar\rho_{1}},$ $\hat{\Phi}=\pi R^{2}B_{ext}/\Phi_{0}$, and
@@
Let’s first consider the simple case where $\mathcal{O}(p^{4})$ terms in Eq. (1) vanish ($\gamma^{-1}=0$) so that $I_{s}=I_{c}\sin\tilde{\phi}$. Here we have an anomalous phase $\phi_{0}$ where the ChNt-JJ has the minimum of its free energy at $\phi_{0}\neq 0$ and $I_{s}(-\phi)\neq-I_{s}(\phi)$ (allowed under broken time-reversal symmetry). In terms of the anisotropy of the nanotube, $\phi_{0}\propto\sin 2\theta/m_{1},$ and in geometry and magnetic field $\phi_{0}\propto B_{ext}A_{n}$ where $A_{n}$ is the surface area of the nanotube in the junction. The latter relationship implies $\phi_{0}$ is linear in the junction length $L$, similar to short Rashba junctions 36. The presence of an anomalous phase here is due to a purely orbital mechanism 38 rather than the more common spin-orbit mechanism 36, 39.
@@
Prior work on a Rashba nanowire JJ with a magnetic field perpendicular to the current flow showed that the anomalous phase due to the spin-orbit interaction 40 is related to a phase shift $\varphi_{0}\propto\Delta v_{F,\sigma}L$ in the Andreev bound state spectrum, where $\Delta v_{F,\sigma}$ is the difference in Fermi velocities of the two spin channels in the junction. Thus, the anomalous phases in Rashba JJs and ChNt-JJs are directly related to broken chiral symmetry in the condensate velocity. This suggests an anomalous phase can arise when the condensate velocity is non-reciprocal, and the JDE develops when the non-reciprocity in the condensate velocity cannot be gauged away in the condensate wavefunction 41.
@@
Figure 2: (a) $\eta$ versus $\theta$ in polar coordinates for $\hat{\Phi}=0.5$ and $L/R=10$. (b) $\eta$ versus $I_{0}/I_{c}$. $\eta$ versus $\theta$ for $\hat{\Phi}=0,~{}0.1,...,~{}0.5$ with (c) $L/R=10$ and (d) $L/R=1$. Here we used $m_{1}/m_{0}=5$, $\kappa_{1}m_{1}/R^{2}=50$, and $\lambda m_{1}/R^{2}=10$.
@@
Returning to the full solution in Eq. (9), we observe the CPR takes an unconventional bipartite form $I_{s}(\phi)=\tilde{I}_{s}(\phi)+I_{0}$ where $I_{0}$ is independent of $\phi$. The phase-independent term $I_{0}$ represents a persistent current in the chiral tube associated with the supercurrent flowing around the tube due to the Little-Parks effect. We have $I_{0}\propto B_{ext}\sin(2\theta)$ so that the current is non-zero only when the tube is chiral and an external magnetic field is applied. While a magnetic field-induced supercurrent around the tube (i.e. $J_{y}\neq 0$) is conventionally expected, it is surprising that this persistent current contributes to the diode effect directly. The picture for the persistent current in this case is depicted in Fig. 1(d), where the current flows along a helical path around the chiral tube. To preserve the requirement that persistent currents form closed loops in equilibrium, the junction will self-tune to the same anomalous phase ($\phi=\phi_{0}$) so that no net supercurrent flows along the $x$-direction.
[…]

### v2

# Perfect supercurrent diode efficiency in chiral nanotube-based weak links

@@
Perfect supercurrent diode efficiency in chiral nanotube-based weak links
Joseph J. Cuozzo jjcuozzo@utep.edu Department of Physics, The University of Texas at El Paso, El Paso, TX, USA. Materials Physics Department, Sandia National Laboratories, Livermore, CA 94551, USA.    François Léonard Materials Physics Department, Sandia National Laboratories, Livermore, CA 94551, USA.
@@
The supercurrent diode effect (SDE) describes superconducting systems where the magnitude of the superconducting-to-normal state switching current differs for positive and negative current bias. Despite the ubiquity of such diode effects in Josephson devices, the fundamental conditions to observe a diode effect in a Josephson junction and achieve perfect diode efficiency remain unclear. In this work, we analyze the supercurrent diode properties of a chiral nanotube-based weak link within a Ginzburg-Landau theory. We find a diode effect and anomalous phase develop across the junction when a magnetic field is applied parallel to the tube despite the absence of spin-orbit interactions in the system. Unexpectedly, the SDE in the junction is independent of the anomalous phase. Alternatively, we determine a non-reciprocal persistent current that is protected by fluxoid quantization can activate SDE, even in the absence of higher-order pair tunneling processes. We show this new type of SDE can lead to, in principle, a perfect diode efficiency, highlighting how persistent currents can be used to engineer high efficiency supercurrent diodes.
I Introduction
@@
With a growing need for low-power fast electronics with low operational temperatures, a rapid development of non-reciprocal superconducting devices has occurred recently 1, 2. Non-reciprocal effects have been intensely investigated in bulk superconductors 3, 4, 5, 6, 7, 8, 9 as well as in Josephson junctions 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27 (JJs)– two superconducting electrodes weakly coupled by a tunneling barrier, such as a normal metal or insulator. In JJs, the Josephson diode effect (JDE) manifests as a difference in magnitude of positive and negative threshold currents where the device switches from a superconducting to a dissipative state, shown schematically in Fig. 1(a). Despite the effect having been known decades ago in devices with geometric inhomogeneities, the subject has received renewed interest based on superconducting diode effects reported in homogeneous devices where non-reciprocity arises from microscopic interactions. In these systems the diode effect, particularly when no magnetic field is applied, has supported the discovery of exotic states, such as time-reversal symmetry broken superconducting states 28, 29. Other zero-field supercurrent diodes include JJs made of iron-based superconductors 30, JJs of twisted bilayer graphene 31, twisted trilayer graphene 32, obstructed atomic insulator JJs 33, strained PbTaSe2 34, and multiferroic JJs 35. Open questions linger about the nature of the SDE in some of these systems and others, which calls for further theoretical modeling to address the fundamental limits of the SDE.
@@
Figure 1: Chiral nanotube-based supercurrent diode: (a) Schematic of voltage-current curve of a supercurrent diode having a negative diode polarity. (b) Cartoon of the ChNt-$J$s and the helical persistent current inducing the Josephson diode effect. (c) Schematic of $\mathcal{C}_{2}$-symmetric Fermi surface. (d) Illustration of chiral nanotube on a square lattice.
@@
Theoretical descriptions of the JDE often heuristically focus on symmetry arguments, namely, broken inversion and time-reversal symmetries 36, 37, 38, 1. Assuming these symmetries are broken, one may consider a common minimal expression for the current-phase relationship (CPR) 2: $I_{s}(\phi)=a\sin(\phi)+b\cos(\phi)+c\sin(2\phi)+d\cos(2\phi).$ Here $\phi$ is the phase across the Josephson junction, and $a,b,c$ and $d$ are real-valued constants describing the weights describing phase-coherent Cooper pair tunneling ($a$ and $b$) and pair co-tunneling ($c$ and $d$) supercurrent channels. The parameters $b$ and $d$ are associated with broken time-reversal symmetry (TRS) in the junction. When all four constants are non-zero and treated as independent parameters, then ubiquitous combinations of $(a,b,c,d)$ result in $I_{c+}\neq|I_{c-}|$, where $I_{c+}=\max\left(I_{s}\right)$ and $I_{c-}=\min\left(I_{s}\right),$ and the Josephson diode effect is realized (see Fig. 1(a)) with an efficiency $\eta=\frac{I_{c+}+I_{c-}}{I_{c+}-I_{c-}}$. Using the minimal form of $I_{s}(\phi)$, it is apparently not possible to have a perfectly efficient supercurrent diode where either $I_{c+}$ or $I_{c-}$ are zero. While ideal diode operation has been identified in the ac limit 39, 40, 41, there remains a question in the dc case: is it possible, in principle, to have a perfectly efficient Josephson diode operating in the dc limit? We answer this question in the affirmative by considering chiral nanotube-based weak link (ChNt-WLs) which serves as a model for short quasi-1D JJs.
@@
In this work, we present a phenomenological Ginzburg-Landau (GL) theory for a ChNt-WL with a magnetic field applied along the tube, see Fig. 1(b). We consider an anisotropic free energy functional that obeys inversion symmetry and choose periodic boundary conditions along non-high symmetry axes to define a chiral nanotube, see Fig. 1(c-d). We first show by numerically solving the GL equations that a ChNt-WL can have a near-perfect SDE. We then analyze the system in the short radius limit where the CPR can be treated analytically. We show that the ChNt-WL develops an anomalous phase despite the absence of spin-orbit coupling. The link between the anomalous phase and JDE has been investigated in JJs that break inversion symmetry 42, but the existence and possible origin of an anomalous phase in ChNt-WLs has not been identified until now. We also find a diode effect in the absence of pair co-tunneling, due to a non-reciprocal persistent current which is protected by fluxoid quantization in the ChNt-WL. In this case, a phase-independent persistent current allows us to place an upper bound on the diode efficiency demonstrating how perfect efficiency can be achieved in ChNt-WLs.
II Model
@@
We model a superconducting ChNt with higher-order terms in the free energy functional 9:
@@
Here $m_{0}<m_{1},m_{2}$, $\zeta_{0},\zeta_{1},\zeta_{2}>0$, $\alpha\propto(T-T_{c})$ and $\beta$ are the usual GL coefficients, and $\mathbf{p}=-i\hbar\left(\nabla-i\frac{2e}{\hbar c}\mathbf{A}\right)$ is the momentum operator. The atomic lattice of the system is defined on a 2D sheet $\Omega\in\mathbb{R}^{2}$. Here the kinetic energy contribution to $F$ has inversion symmetry and $\mathcal{C}_{2}$ rotational symmetry when $1/m_{1}\neq 0.$ This corresponds to an elliptical Fermi surface, see Fig. 1(c). The circumferential vector $\mathbf{C}=2\pi R(-\sin\theta,\cos\theta),$ where $R$ is the radius of the nanotube, defines the periodic boundary conditions to apply to the nanotube, see Fig. 1(d). To simplify our analysis, we rotate our real space basis $\mathbf{r}=R(\theta)\mathbf{r}_{0}$ by the chiral angle $\theta$ where $R(\theta)$ is the 2x2 rotation matrix. The Ginzburg-Landau equations in this case are (see Supplementary Information for details)
@@
$\left[\alpha+\beta{\left|\psi\right|}^{2}-\left(\rho_{1}p_{x}^{2}+\rho_{2}p_{y}^{2}+2\rho_{3}p_{x}p_{y}\right)+\rho_{4}\left(p_{x}^{4}+p_{y}^{4}\right)+\left(\rho_{6}+2\rho_{7}\right)p_{x}^{2}p_{y}^{2}+2\rho_{8}\left(p_{x}^{3}p_{y}-p_{x}p_{y}^{3}\right)\right]\psi=0,$ (3)
@@
To establish the conditions for the SDE, we consider an external magnetic field along the nanotube $\mathbf{B}=B_{ext}\mathbf{x}$ which breaks time-reversal symmetry. Periodic boundary conditions lead to $p_{y}$ being a good quantum number $p_{y}^{(n)}=\hbar(\pi R^{2}B_{ext}/\Phi_{0}+n)/R=\hbar(\hat{\Phi}+n)/R$ for some $n\in\mathbb{Z}$ and where $\Phi_{0}$ is the flux quantum and $\hat{\Phi}=\Phi/\Phi_{0}$. From here, to calculate the supercurrent in a Josephson junction, we need to apply appropriate boundary conditions to the solution to Eq. (3) in the weak link. The simplest model for a narrow constriction uses rigid boundary conditions 43
$\psi(x\leq 0)=\psi_{\infty},\quad\psi(x\geq L)=\psi_{\infty}e^{i\phi},$ (4)
where the region defining the junction is $0\leq x\leq L$, $\phi$ is the phase difference across the junction, and $\psi_{\infty}=\sqrt{-\alpha/\beta}.$ The supercurrent density calculated in the weak link region describes the flow of Cooper pairs, both with and without $\phi$-sensitivity. The $\phi$-dependent contribution to the supercurrent density describes phase-coherent transport of Cooper pairs across the junction, which is the Josephson current. The $\phi$-independent term is a persistent current. This approach within GL theory qualitatively captures the current-phase relationship of Josephson junction 44 and has been successfully used before to describe Rashba JJs under an applied magnetic field 43. A more rigorous treatment of the boundary conditions with a self-consistent treatment of $\psi$ for the junction can be made and may be important in 2D and 3D geometries 45, but for our quasi-1D system rigid boundary conditions provide sufficient qualitative accuracy 44. In superconductor-quantum dot-superconductor JJs, Coulomb interactions have been predicted to influence the CPR and induce a zero field SDE 26, 27. In our model, we assume the Josephson energy of the weak link is greater than the charging energy so that the phase difference across the junction can be treated classically and effects of Coulomb interactions are negligible. Here we also assume $T\lesssim T_{c}$, and take the short junction limit $L\ll\xi$ where $\xi=\sqrt{\hbar^{2}/(2m_{0}|\alpha|)}$ is the GL superconducting coherence length.
Figure 2: Approaching perfect supercurrent diode efficiency: (a) $\eta$ versus $\hat{\Phi}$ for various ratios of $L/R$, $m_{1}/m_{0}=100$, and $\theta=0.1\pi$. (b) $\eta$ versus $\theta$ for $\hat{\Phi}=0.1,~0.2,~0.3,~0.4$, $m_{1}/m_{0}=100$, and $L/R=10^{-3}$. (c) $\eta$ versus $m_{1}$ for various ratios of $L/R$ and with $\hat{\Phi}=0.1$ and $\theta=0.1\pi$. (d) CPR of the junction for various ratios of $L/R$ with $\hat{\Phi}=0.1$, $m_{1}/m_{0}=100$, and $\theta=0.1\pi$ using the normalization $I_{c}=\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}$. We numerically solved Eq. (2-3) as described in the main text with $\frac{\hbar^{2}}{m_{0}\zeta_{0}L^{2}}=2\times 10^{-3}$, $\frac{\hbar^{2}m_{0}}{m_{1}^{2}\zeta_{0}L^{2}}=\frac{\hbar^{2}m_{0}}{m_{1}^{2}\zeta_{0}L^{2}}=2\times 10^{-4}$, and $\frac{\alpha m_{0}L^{2}}{\hbar^{2}}=10^{-4}$.
III Results
III.1 Numerical solution
We solve Eq. (3) with separation of variables using the order parameter
$\psi_{n}(x,y)$ $=\psi_{x}(x)e^{ip_{y}^{(n)}y/\hbar},$ (5)
for some complex-valued function $\psi_{x}$. In the short junction limit we can ignore the cubic $|\psi|^{2}\psi$ term in Eq. (3) and reduce the equation to a linear fourth-order differential equation for $\psi_{x}$. In the following, we will also ignore $d^{3}\psi_{x}/dx^{3}$ and $d^{4}\psi_{x}/dx^{4}$ terms in solving the differential equation for two reasons: the coefficients of these terms are negligibly small in all simulations we present, and $d^{m}\psi_{x}/dx^{m}$ for $m>2$ are small in the short junction limit. We have numerically verified the numerical solutions we find are consistent with these approximations. In order to calculate the diode efficiency, we numerically calculate $\psi_{x}(x)$ to find $\psi_{n}(x,y)$ and then, following the application of boundary conditions, we calculate the supercurrent density $J_{x}(\phi,n)$ using Eq. (2). Integrating $J_{x}(\phi,n)$ over the junction region, we calculate the CPR $I_{s}(\phi,n_{eq})$ by minimizing the free energy in Eq. (1) with respect to $n$ and $\phi$ to find the value of $n$ at equilibrium, $n_{eq}$, and then treat $\phi$ as an externally tunable parameter.
[…]
Figure 3: Non-reciprocal persistent supercurrent: Phase-integrated CPR over a $2\pi$ cycle versus (a) $\hat{\Phi}$ for various ratios of $L/R$, and (b) $\theta$ for $\hat{\Phi}=0.1,~0.2,~0.3,~0.4$ using the normalization $I_{c}=\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}$. Here we used numerically solved Eq. (2-3) as described in the main text with $m_{1}/m_{0}=100$, $\frac{\hbar^{2}}{m_{0}\zeta_{0}L^{2}}=2\times 10^{-3}$, $\frac{\hbar^{2}m_{0}}{m_{1}^{2}\zeta_{0}L^{2}}=\frac{\hbar^{2}m_{0}}{m_{1}^{2}\zeta_{0}L^{2}}=2\times 10^{-4}$, and $\frac{\alpha m_{0}L^{2}}{\hbar^{2}}=10^{-4}$.
In Fig. 2(b) we present $\eta$ of the ChNt-WL as a function of the chiral angle $\theta$ of the nanotube. Here we see $\eta$ is $\pi$-periodic in $\theta$, reflecting the $C_{2}$ rotational symmetry of the GL free energy functional, Eq. (3). The suppression of $\theta$ at $\pi/2$ and $3\pi/2$ arises from the suppression of $\rho_{3}\propto\sin 2\theta$ in the superfluid stiffness tensor. Figure 2(c) shows $\eta$ versus $m_{1}$ (which controls the superfluid stiffness anisotropy to leading order) for $\hat{\Phi}=0.1$. In the limit $m_{1}\gg m_{0}$, the superfluid stiffness becomes nearly isotropic and $\eta\rightarrow 0$. In the opposite limit $m_{1}\lesssim m_{0}$, $\eta$ approaches perfect diode efficiency regardless of nanotube radius. Importantly, this behavior of $\eta$ occurs despite the small values of $\rho_{4},~\rho_{6},~\rho_{7}$ and $\rho_{8}$ used in the calculations. This is significant since in the absence of higher-order momentum terms in the free energy, the diode effect is completely suppressed. Here we find that even with small contributions from higher-order terms, large anisotropy in lowest-order kinetic terms can induce a large SDE. In Fig. 2(d), we present the CPR of the ChNt-WL for various $L/R$ ratios and with $\hat{\Phi}=0.1$ and $\theta=0.1\pi$. As $L/R$ increases, we observe the combination of a phase shift and current offset in $I_{s}$ such that a diode effect emerges but the phase that minimizes the free energy of the ChNt-JJ remains near zero. While phase shifts in the CPR of JJs when both time-reversal and inversion symmetries are broken have been reported before 43, 46, the current offset indicates there is an unexpected persistent current in the weak link. To investigate this further, in Fig. 3 we calculate the phase-integrated CPR $\int_{0}^{2\pi}I_{s}(\phi)\;d\phi$ which is non-zero when a persistent current is present in the weak link. In Fig. 3(a), we see the phase-integrated CPR is non-zero when $\eta$ is non-zero (c.f. Fig. 2(a)) and is linearly proportional to $\hat{\Phi}$. Similarly, we see in Fig. 3(b) that the persistent current is roughly proportional to $\eta$. This suggests the persistent current is crucial to achieving a near-perfect supercurrent diode efficiency in a ChNt-WL. In the following section, we analytically analyze the origin of the non-reciprocal persistent current in the weak link.
III.2 Small nanotube diameter limit
To develop a clear picture for the SDE in ChNt-WLs, we consider the simpler case of a small diameter nanotube where we restrict ourselves to the lowest sub-band for simplicity, and henceforth set $p_{y}^{(n)}=p_{y}^{(0)}$. Then the order parameter is
@@ III.2 Small nanotube diameter limit
for some complex-valued function $\psi_{x}$. We can solve Eq. (3) by linearizing in the standard way, ignoring higher-order gradients along the axial direction as before, and find
$\psi_{x}(x)$ $\approx\psi_{\infty}L^{-1}e^{iax}\left(L-x+xe^{i(\phi-\phi_{0})}\right),$ (7)
where $\phi_{0}=-\frac{\rho_{3}p_{y}^{(0)}L}{\hbar\rho_{1}}$. Then the current-phase relationship of the ChNt-WL is calculated by solving for $J_{x}$ in Eq. (2) which, due to the anisotropic superfluid stiffness, requires both components of $\bf j$:
$\bf{j}$ $=\begin{pmatrix}2i\frac{\psi_{\infty}^{2}}{L}\sin\left(\phi-\phi_{0}\right)+2i\frac{\psi_{0}^{2}\phi_{0}}{L}{\left|L-x+xe^{i(\phi-\phi_{0})}\right|}^{2}\\ \frac{2i\psi_{0}^{2}p_{y}^{(0)}}{\hbar}{\left|L-x+xe^{i(\phi-\phi_{0})}\right|}^{2}\end{pmatrix}$. (8)
@@
[…]
Here $I_{c}=\frac{4e\hbar\rho_{1}\psi_{\infty}^{2}A_{\perp}}{L}$, $\tilde{\phi}=\phi-\phi_{0}$, $\hat{\Phi}=\pi R^{2}B_{ext}/\Phi_{0}$, and
@@
Let’s first consider the simple case where $\mathcal{O}(p^{4})$ terms in Eq. (1) vanish ($\gamma^{-1}=0$) so that $I_{s}=I_{c}\sin\tilde{\phi}$. Here we have an anomalous phase $\phi_{0}$ where the ChNt-WL has the minimum of its free energy at $\phi_{0}\neq 0$ and $I_{s}(-\phi)\neq-I_{s}(\phi)$ (allowed under broken time-reversal symmetry). In terms of the anisotropy of the nanotube, $\phi_{0}\propto\sin 2\theta/m_{1},$ and in geometry and magnetic field $\phi_{0}\propto B_{ext}A_{n}$ where $A_{n}$ is the surface area of the nanotube in the junction. The latter relationship implies $\phi_{0}$ is linear in the junction length $L$, similar to short Rashba junctions 43. The presence of an anomalous phase here is due to a purely orbital mechanism 47 rather than the more common spin-orbit mechanism 43, 46.
@@
Prior work on a Rashba nanowire JJ with a magnetic field perpendicular to the current flow showed that the anomalous phase due to the spin-orbit interaction 48 is related to a phase shift $\varphi_{0}\propto\Delta v_{F,\sigma}L$ in the Andreev bound state spectrum, where $\Delta v_{F,\sigma}$ is the difference in Fermi velocities of the two spin channels in the junction. Thus, the anomalous phases in Rashba JJs and ChNt-WLs are directly related to broken chiral symmetry in the condensate velocity. This suggests an anomalous phase can arise when the condensate velocity is non-reciprocal, and the JDE develops when the non-reciprocity in the condensate velocity cannot be gauged away in the condensate wavefunction 49.
@@
Figure 4: Optimizing diode efficiency: (a) $|\eta|$ versus $\theta$ in polar coordinates for $\hat{\Phi}=0.5$ and $L/R=10$. (b) $\eta$ versus $I_{0}/I_{c}$. $\eta$ versus $\theta$ for $\hat{\Phi}=0,~0.1,...,~0.5$ with (c) $L/R=10$ and (d) $L/R=1$. Here we used $m_{1}/m_{0}=5$, $\kappa_{1}m_{1}/R^{2}=50$, and $\lambda m_{1}/R^{2}=10$ in Eq. (13).
@@
Returning to the full solution in Eq. (10), we observe the CPR takes an unconventional bipartite form $I_{s}(\phi)=\tilde{I}_{s}(\phi)+I_{0}$ where $I_{0}$ is independent of $\phi$. The phase-independent term $I_{0}$ represents a persistent current in the chiral tube associated with the supercurrent flowing around the tube due to the Little-Parks effect. We have $I_{0}\propto B_{ext}\sin(2\theta)$ so that the current is non-zero only when the tube is chiral and an external magnetic field is applied. While a magnetic field-induced supercurrent around the tube (i.e. $J_{y}\neq 0$) is conventionally expected, it is surprising that this persistent current contributes to the diode effect directly. The picture for the persistent current in this case is depicted in Fig. 1(b), where the current flows along a helical path around the chiral tube. To preserve the requirement that persistent currents form closed loops in equilibrium, the junction will self-tune to the same anomalous phase ($\phi=\phi_{0}$) so that no net supercurrent flows along the $x$-direction.
We can compare the persistent current in Eq. (10) to the phase-independent supercurrent in the CPR of a junction with Meissner screening inducing non-reciprocity 19. In Ref. 19, the Meissner effect generates a persistent current due to spectral flow of Andreev bound states that leads to $\int_{0}^{2\pi}I_{s}d\phi=0$, but in our work we find $\int_{0}^{2\pi}I_{s}d\phi\neq 0$. The screening current in Ref. 19 is found to induce a diode effect only when higher harmonics enter the CPR i.e. the diode effect vanishes in the CPR to second order in the tunneling amplitude. This suggests the persistent current generated by the Meissner effect in the superconducting electrodes enters higher-order pair tunneling channels across the junction, leading to an interference between channels that causes a diode effect. In our case, the persistent current leading to a diode effect is not dependent on pair channel interference or any phase-coherent Josephson tunneling process. The persistent current is associated with $\mathcal{O}(p^{4})$ terms in Eq. (1), and we found replacing $\mathcal{O}(p^{4})$ terms with terms describing pair co-tunneling does not necessarily result in a diode effect (see the SI). Furthermore, the persistent current is protected by fluxoid quantization so that when the order parameter $\psi$ is non-zero in the ChNt-WL, the non-reciprocal component of the persistent current generally flows across the junction. A similar persistent current is also found in asymmetric SQUIDs 39. There, fluxoid quantization dictates the magnetic properties of the device and screening introduces a non-reciprocal persistent current contributing to the supercurrent diode effect, but pair co-tunneling is also found to be necessary for the diode effect 39. Thus, ChNt-WLs exhibit a new type of SDE with persistent currents where pair co-tunneling is unnecessary.